you have sold 36 units of product Y. you are currently at 80% of you goal. how many more units need to be sold before you reach your goal?
Answer by lwsshak3(11628)

you have sold 36 units of product Y. you are currently at 80% of you goal. how many more units need to be sold before you reach your goal?
**
let x=total number of units need to be sold to reach your goal
36/x=.8
x=36/.8=45
ans:
additional units need to be sold to reach your goal=45-36=9